OpenStudy (lgbasallote):

\[\large \int \frac{1}{\sqrt{1 - t^4}}dt\]

OpenStudy (dumbcow):

initial guess... t = sqrt(sin u)

OpenStudy (anonymous):

That makes dt=.5cos u/sqrt(sin u), which might be a bit of a pain... but that was my initial guess too.

OpenStudy (lgbasallote):

for the record...i do not know how to solve this...

OpenStudy (anonymous):

Oh, 1-t^4=(1-t^2)(1+t^2), so just do t=sin u

OpenStudy (anonymous):

It will still be awkward though because of the 1+t^2 part. And if you do t=tan u then the 1-t^2 part will be awkward. hmmmm

OpenStudy (anonymous):

suppose t@2 is= u

OpenStudy (dumbcow):

that gives \[\frac{1}{2}\int\limits_{?}^{?}\frac{1}{\sqrt{u} \sqrt{1-u^{2}}} du\]

OpenStudy (dumbcow):

you can all give up now... http://www.wolframalpha.com/input/?i=integrate+1%2Fsqrt%281-t^4%29+dt hmm elliptical integrals

OpenStudy (anonymous):

Haha thank you for checking that, I would have been working on this for another 15-30 minutes :P

OpenStudy (lgbasallote):

courtesy of @apoorvk lol

OpenStudy (apoorvk):

Lol @lgbasallote